# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v2

@@ Abstract
We investigate $\phi^{2n+1}$ deformations of the generalized free theory in the $\epsilon$ expansion, where the canonical kinetic term is generalized to a higher-derivative version. For $n=1$, we use the conformal multiplet recombination method to determine the leading anomalous dimensions of the fundamental scalar operator $\phi$ and the bilinear composite operators $\mathcal{J}$. Then we extend the $n=1$ analysis to the Potts model with $S_{N+1}$ symmetry and its higher-derivative generalization, in which $\phi$ is promoted to an $N$-component field. We further examine the Chew-Frautschi plots and their $N$ dependence. However, for each integer $n>1$, the leading anomalous dimensions of $\phi$ and $\mathcal{J}$ are not fully determined and contain one unconstrained constant, which in the canonical cases can be fixed by the results from the traditional diagrammatic method. In all cases, we verify that the multiplet-recombination results are consistent with crossing symmetry using the analytic bootstrap methods.
@@ 1 Introduction
For a scalar field theory, the canonical kinetic term is typically proportional to $\partial_{\mu}\phi\partial^{\mu}\phi$, [^1] Summation over repeated indices is implied. with $\phi$ denoting the fundamental scalar. This kinetic term can be rewritten as $\phi\,\Box\phi$ up to a total derivative which we discard. The expression $\phi\,\Box\phi$ can be generalized by employing a higher-order Laplacian. In this paper, we use the word “generalized” to indicate a free theory with a higher-derivative kinetic term and the interacting deformations. The action of the generalized free theory reads
@@
where $d$ is the spacetime dimension and $k$ is a positive integer (see Brust:2016gjy for more details). One can introduce interaction terms to (1) and investigate the infrared (IR) fixed points of the renormalization group (RG) flow (see for example David:1992vv ; Gracey:2017erc ; Safari:2017irw ; Safari:2017tgs ). We assume that the IR fixed points possess conformal symmetry, allowing for the use of conformal field theory (CFT) methods to study the critical properties.
@@
The first instance of these $\mathbb{Z}_{2}$-odd theories corresponds to $n=1$, i.e., the generalized $\phi^{3}$ theory. We also refer to this theory as the generalized Yang-Lee edge singularity, since the canonical $\phi^{3}$ theory describes the Yang-Lee edge singularity Fisher:1978pf .
Before moving on to higher $n$, we also study the generalized $\phi^{3}$ theory with $S_{N+1}$ global symmetry, [^3] One may also introduce an auxiliary field to define an O($N$) symmetric theory in $d=6-\epsilon$ dimensions Fei:2014yja . At large $N$, it is conjectured that the IR CFT at $d=5$ is dual to the Vasiliev higher-spin theory on $\text{AdS}_{6}$. where we promote $\phi$ to an $N$-component field $\vec{\phi}$. The symmetric group $S_{N+1}$ is the group of all permutations of $N+1$ objects. It can also be viewed as the symmetry group of an $N$-simplex. For example, $S_{3}$ is isomorphic to the symmetry group of an equilateral triangle, and this group consists of 6 elements: the identity, rotations by $32\pi/$ 3 and $34\pi/$ 3 about the center, as well as reflections about the three symmetry axes of the equilateral triangle. The mismatch between the number of components of $\phi$ and the degree of the symmetric group originates from the lattice theory described in the next paragraph. In the generalized $S_{N+1}$-symmetric $\phi^{3}$ theory, the $\mathbb{Z}_{2}$ symmetry under the reflection $\vec{\phi}\rightarrow-\vec{\phi}$ is absent, as in the Yang-Lee case. [^4] The symmetric group $S_{N+1}$ contains $\mathbb{Z}_{2}$ subgroups, but the elements of these subgroups do not act by $\vec{\phi}\rightarrow\pm\vec{\phi}$. See appendix C. At $k=1$, this theory describes the $(N+1)$-state Potts model, so we also refer to the generalized $S_{N+1}$-symmetric $\phi^{3}$ theory as the generalized Potts model.
@@
After discussing the $n=1$ case, we proceed to consider theories with greater integer $n$ [^9] For studies of the canonical $n=2$ theory with $S_{N+1}$ symmetry, see Codello:2018nbe ; Codello:2020mnt . We only consider the single-field case for the $n>1$ theory in this paper. For $n=2$, the interaction is associated with $\phi^{5}$, and the corresponding CFT is expected to be a multicritical generalization of the Yang-Lee CFT. More broadly, we consider the generalized $\phi^{2n+1}$ theory with $n>1$. A notable feature is that, at leading order in the $\epsilon$ expansion, a family of data is consistent with conformal symmetry, as far as our analysis is concerned. This means that a large set of leading anomalous dimensions is determined up to one parameter, in contrast to the $n=1$ case where the anomalous dimensions are fixed. As the $\phi^{5}$ theory, the $\phi^{2n+1}$ theories with higher $n$ are also associated with multicritical Yang-Lee models vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib , and they provide a series of universality classes containing multicritical points with $\mathbb{Z}_{2}$-odd interactions. See also Zambelli:2016cbw ; Gracey:2017okb ; Codello:2017qek ; Codello:2017epp for related studies. This concludes our discussion of the theories that will be investigated in this paper.
@@
where the indices are suppressed for simplicity. The partially conserved currents do not vanish when we act with fewer derivatives. These operators are related to higher spin symmetry of the free theory. One can construct conserved charges from the partially conserved currents, generalizing the $k=1$ higher spin algebra. We refer to Brust:2016gjy for more details. For $k>1$, the interacting versions of the partially conserved currents play a significant role in the analytic bootstrap, as their twists are lower than those of the usual conserved currents. In the lightcone limit, the interacting counterparts of the partially conserved currents are more important than the Regge trajectories associated with the conserved currents.
@@
To derive the anomalous dimensions of the bilinear operators, we employ the conformal multiplet recombination method initiated in Rychkov:2015naa , which considered the canonical $\phi^{4}$ theory. In Rychkov:2015naa , it was further suggested that the results can be extended to the canonical $\phi^{3}$ theory at $d=6-\epsilon$.  [^10] According to Rychkov:2015naa ; Gliozzi:2016ysv ; Gliozzi:2017gzh , Yu Nakayama has considered the multiplet recombination for the $\epsilon=6-d$ expansion of the $\phi^{3}$ theory in his unpublished notes. A key difference is the absence of $\mathbb{Z}_{2}$ symmetry in the $\phi^{3}$ theory, in contrast to the $\phi^{4}$ case. Since the procedure in Rychkov:2015naa relies on the $\mathbb{Z}_{2}$ symmetry, its generalization to the $\mathbb{Z}_{2}$-odd cases does not seem straightforward. A similar approach based on the equation of motion was proposed to study the $\phi^{3}$ theory Nii:2016lpa . By applying the equation of motion to correlation functions, the constraints on the anomalous dimensions can be derived more readily. In this paper, we revisit the $\phi^{2n+1}$ theory from the multiplet recombination perspective and extend the results to some spinning operators. As a consistency check, we also verify that the results for the anomalous dimensions and OPE coefficients are consistent with crossing symmetry using the analytic bootstrap.
@@
The multiplet recombination asserts that free multiplets recombine into long multiplets at interacting IR fixed points. For instance, at $d=4-\epsilon$, the Wilson-Fisher multiplet $\{\phi\}$ is a long multiplet, into which the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{3}_{\text{f}}\}$ recombine. Here the subscript f implies the free theory. In the Gaussian limit, the descendant $\Box\phi$ is expected to become null and decouple, but a singular change in the normalization can lead to a physical operator, which is identified with the free theory primary operator $\phi^{3}_{\text{f}}$. In other words, the Wilson-Fisher counterpart of $\phi^{3}_{\text{f}}$ is identified as a descendant $\Box\phi$ up to a normalization factor  [^11] All physical operators in the free theory should have counterparts in the interacting theory.:
@@
Despite the resemblance to the equation of motion $\Box\phi\propto\epsilon\,\phi^{3}$, the multiplet recombination is conceptually different, as it can be formulated purely in the CFT language and does not require a Lagrangian description. The multiplet recombination above generalizes to various cases. As in Guo:2023qtt , we follow a suggestion of Gliozzi:2017gzh and we apply the multiplet recombination method to the $\mathbb{Z}_{2}$-odd deformations of the free higher-derivative CFTs. See Basu:2015gpa ; Yamaguchi:2016pbj ; Ghosh:2015opa ; Raju:2015fza ; Gliozzi:2016ysv ; Roumpedakis:2016qcg ; Soderberg:2017oaa ; Behan:2017emf ; Gliozzi:2017hni ; Gliozzi:2017gzh ; Nishioka:2022odm ; Nishioka:2022qmj ; Antunes:2022vtb ; Guo:2023qtt for applications of the multiplet recombination to numerous CFTs, and Safari:2017irw ; Nii:2016lpa ; Hasegawa:2016piv ; Hasegawa:2018yqg ; Skvortsov:2015pea ; Giombi:2016hkj ; Giombi:2017rhm ; Giombi:2016zwa ; Codello:2017qek ; Codello:2018nbe ; Antipin:2019vdg ; Vacca:2019rsh ; Giombi:2020rmc ; Giombi:2020xah ; Dey:2020jlc ; Giombi:2021cnr ; Safari:2021ocb ; Zhou:2022pah ; Bissi:2022bgu ; Herzog:2022jlx ; Giombi:2022vnz ; SoderbergRousu:2023pbe for applications with more emphasis on the equations of motion.
This paper is organized as follows. In section 2, we consider the generalized Yang-Lee edge singularity. We first employ the multiplet recombination to compute the anomalous dimension of $\phi$ and bilinear operators $\mathcal{J}$, then we use the analytic bootstrap to verify that the results are consistent with crossing symmetry. In section 3, we carry out a similar analysis, obtaining the anomalous dimensions in the generalized Potts model, and we check that the results are consistent with crossing symmetry as well. Section 4 focuses on the generalized $\phi^{2n+1}$ theory with $n>1$, where we only consider the single-field case. We determine the anomalous dimensions up to a free parameter, then we verify the consistency between the multiplet-recombination results and crossing symmetry. We summarize the main results in table 1. In appendix A, we work out the action of Laplacians on the scalar three-point function. In appendix B, we verify, in all three cases, that the multiplet-recombination results are consistent with crossing symmetry at subleading twist. We briefly review some representation theory of the symmetric group in appendix C. More details about the tensor structures in the Potts model can be found in appendix D. We compute in appendix E the ratios of OPE coefficients relevant to the multiplet recombination. In appendix F, we derive an OPE coefficient used in the analytic bootstrap. We discuss in appendix G the $N\rightarrow 0$ limit of the generalized Potts model in more detail, especially the logarithmic structures. Appendix H discusses some special limits in the generalized O($N$)-symmetric $\phi^{2n}$ theory.
@@ 2 Generalized Yang-Lee edge singularity
Let us first consider the generalized Yang-Lee CFT. At $d=d_{\text{u}}-\epsilon$, the action reads [^12] We do not need to consider derivative interactions. The reason is that a derivative interaction should be constructed by fewer $\phi$’s, namely two $\phi$’s, but a derivative term with two $\phi$’s is just the kinetic term, up to a total derivative.
$S=\int\mathrm{d}^{d}x\left(\phi\,\Box^{k}\phi+g\mu^{\epsilon/2}\phi^{3}\right)\,,$ (5)
@@ 2.1 Multiplet recombination
where $\alpha=\alpha(\epsilon)$ vanishes in the Gaussian limit: $\lim_{\epsilon\rightarrow 0}\alpha=0$. The above equation means that we identify the descendant $\alpha^{-1}\Box^{k}\phi$ as the deformed version of $\phi^{2}_{\text{f}}$, and the normalization $\alpha^{-1}$ is introduced to ensure a finite Gaussian limit. In other words, the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{2}_{\text{f}}\}$ recombine into a long multiplet $\{\phi\}$.
@@
Our focus will be on the two-point and three-point functions. We study the matching conditions for the scalar correlators first. As we will see below, these conditions determine the anomalous dimension of $\phi$ and the normalization $\alpha$, as well as the OPE coefficient associated with $\phi\times\phi\sim\phi$, at leading order in the $\epsilon$ expansion. Then, we consider the matching condition for the spinning correlator involving a primary bilinear operator. [^13] The spin of the bilinear operator can be zero. Using the results for the anomalous dimension of $\phi$ and $\alpha$, we obtain the leading anomalous dimension of the bilinear operator.
@@
On the right-hand side of (7), the free correlator is given by Wick’s theorem. [^14] We assume that the Gaussian limit is smooth, so the normalization in the free theory is also fixed by (9), i.e., $\langle\phi_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\rangle=1/x_{12}^{\smash{2\Delta_{\phi_{\text{f}}}}}$. The functional forms in $x_{12}$ indeed match in (8), since
@@
and substitute $d=6k-\epsilon$ in (8). Omitting the functional forms in $x_{12}$, we are left with
@@
As before, the right-hand side is computed using Wick’s theorem. On the left-hand side, we act on $\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ with the Laplacians. The result is worked out in appendix A. We obtain
@@
where $\lambda_{\phi\phi\phi}$ is the OPE coefficient of $\phi\in\phi\times\phi$. The coefficient consists of a part independent of the scaling dimensions:
$B_{i,j}$ $=\frac{(-1)^{i+j-k}2^{2k}k!}{(k-i)!\,(k-j)!\,(i+j-k)!}\,,$ (16)
and another part given by
$C_{i,j}^{a,b}$ $=\Big{(}\frac{a}{2}\Big{)}_{i}\Big{(}\frac{b}{2}\Big{)}_{j}\left(\frac{a+b}{2}-\frac{d-2}{2}+i+j-k\right)_{2k-i-j}\,.$ (17)
Due to the last Pochhammer symbol, $C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}$ is proportional to the anomalous dimension $\gamma_{\phi}$, except for the $i=j=k$ case where $C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}$ is finite in the Gaussian limit. In other words, the dominant term in the $\epsilon\rightarrow 0$ limit is given by $i=j=k$. Then, one can check that the functional forms match in (15). We obtain a constraint on $\lambda_{\phi\phi\phi}$ and $\alpha$:
@@
indicating that the OPE coefficient $\lambda_{\phi\phi\phi}$ is of order $\alpha$. Another matching condition is
@@
The action of the Laplacians is derived in appendix A. We have
$\lim_{\epsilon\rightarrow 0}\Bigg{\{}\alpha^{-3}$ $\sum^{k}_{\begin{subarray}{c}i_{1},j_{1}=0\\ i_{1}+j_{1}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{2},j_{2}=0\\ i_{2}+j_{2}\geqslant k\end{subarray}}\;\;\sum^{k}_{\begin{subarray}{c}i_{3},j_{3}=0\\ i_{3}+j_{3}\geqslant k\end{subarray}}\Bigg{[}B_{i_{1},j_{1}}C_{i_{1},j_{1}}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}B_{i_{2},j_{2}}C_{i_{2},j_{2}}^{\Delta_{\smash{\phi}}+2i_{1},\Delta_{\smash{\phi}}+2(k-i_{1}-j_{1})}$
×$\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{\Delta_{\smash{\phi}}+2(j_{2}+k-i_{1}-j_{1}),\Delta_{\smash{\phi}}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{\lambda_{\phi\phi\phi}}{x_{12}^{\Delta_{\smash{\phi}}}x_{13}^{\Delta_{\smash{\phi}}}x_{23}^{\Delta_{\smash{\phi}}}}\Bigg{\}}=\frac{8}{x_{12}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}x_{13}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}x_{23}^{\Delta_{\smash{\phi^{2}_{\text{f}}}}}}\,,$ (20)
where $B_{i,j}$ and $C_{i,j}^{a,b}$ are defined in (16) and (17). As before, the functional forms match, leaving us with a constraint on $\gamma_{\phi}$, $\lambda_{\phi\phi\phi}$ and $\alpha$:

### v3

@@ Abstract
We investigate $\phi^{2n+1}$ deformations of the generalized free theory in the $\epsilon$ expansion, where the canonical kinetic term is generalized to a higher-derivative version. For $n=1$, we use the conformal multiplet recombination method to determine the leading anomalous dimensions of the fundamental scalar operator $\phi$ and the bilinear composite operators $\mathcal{J}$. Then we extend the $n=1$ analysis to the Potts model with $S_{N+1}$ symmetry and its higher-derivative generalization, in which $\phi$ is promoted to an $N$-component field. We further examine the Chew-Frautschi plots and their $N$ dependence. However, for each integer $n>1$, the leading anomalous dimensions of $\phi$ and $\mathcal{J}$ are not fully determined and contain one unconstrained constant, which in the canonical cases can be fixed by the results from the diagrammatic method. In all cases, we verify that the multiplet-recombination results are consistent with crossing symmetry using the analytic bootstrap methods.
@@ 1 Introduction
For a scalar field theory, the canonical kinetic term is typically proportional to $\partial_{\mu}\phi\partial^{\mu}\phi$, [^1] Summation over repeated indices is implied. with $\phi$ denoting the fundamental scalar. This kinetic term can be rewritten as $\phi\,\Box\phi$ up to a total derivative, which we discard. The expression $\phi\,\Box\phi$ can be generalized by employing a higher-order Laplacian. In this paper, we use the word “generalized” to indicate a free theory with a higher-derivative kinetic term and the interacting deformations. The action of the generalized free theory reads
@@
where $d$ is the spacetime dimension and $k$ is a positive integer (see Brust:2016gjy for more details). One can introduce interaction terms into (1) and investigate the infrared (IR) fixed points of the renormalization group flow (see, for example David:1992vv ; Gracey:2017erc ; Safari:2017irw ; Safari:2017tgs ). We assume that the IR fixed points possess conformal symmetry, allowing for the use of conformal field theory (CFT) methods to study the critical properties.
@@
The first instance of these $\mathbb{Z}_{2}$-odd theories corresponds to $n=1$, i.e., the generalized $\phi^{3}$ theory. We also refer to this theory as the generalized Yang-Lee edge singularity, [^3] The $k=2$ Yang-Lee theory might be related to the zeros of the partition function of the Lifshitz theory. since the canonical $\phi^{3}$ theory describes the Yang-Lee edge singularity Fisher:1978pf .
Before moving on to higher $n$, we also study the generalized $\phi^{3}$ theory with $S_{N+1}$ global symmetry, [^4] One may also introduce an auxiliary field to define an O($N$) symmetric theory in $d=6-\epsilon$ dimensions Fei:2014yja . At large $N$, it is conjectured that the IR CFT at $d=5$ is dual to the Vasiliev higher-spin theory on $\text{AdS}_{6}$. where we promote $\phi$ to an $N$-component field $\vec{\phi}$. The symmetric group $S_{N+1}$ is the group of all permutations of $N+1$ objects. It can also be viewed as the symmetry group of an $N$-simplex. For example, $S_{3}$ is isomorphic to the symmetry group of an equilateral triangle, and this group consists of six elements: the identity, rotations by $32\pi/$ 3 and $34\pi/$ 3 about the center, and reflections about the three symmetry axes of the equilateral triangle. The mismatch between the number of components of $\phi$ and the degree of the symmetric group originates from the lattice theory described in the next paragraph. In the generalized $S_{N+1}$-symmetric $\phi^{3}$ theory, the $\mathbb{Z}_{2}$ symmetry under the reflection $\vec{\phi}\rightarrow-\vec{\phi}$ is absent, as in the Yang-Lee case. [^5] The symmetric group $S_{N+1}$ contains $\mathbb{Z}_{2}$ subgroups, but the elements of these subgroups do not act by $\vec{\phi}\rightarrow\pm\vec{\phi}$. See appendix C. At $k=1$, this theory describes the $(N+1)$-state Potts model, so we also refer to the generalized $S_{N+1}$-symmetric $\phi^{3}$ theory as the generalized Potts model. [^6] The critical point of the $k=2$ Potts model may be an $S_{N+1}$-symmetric version of an isotropic Lifshitz point.
@@
After discussing the $n=1$ case, we proceed to consider theories with greater integer $n$ [^11] See Codello:2018nbe ; Codello:2020mnt for the studies of the canonical $n=2$ theory with $S_{N+1}$ symmetry. We only consider the single-field case for the $n>1$ theory in this paper. For $n=2$, the interaction is associated with $\phi^{5}$, and the corresponding CFT is expected to be a multicritical generalization of the Yang-Lee CFT. More broadly, we consider the generalized $\phi^{2n+1}$ theory with $n>1$. A notable feature is that, at leading order in the $\epsilon$ expansion, a family of data is consistent with conformal symmetry, as far as our analysis is concerned. This means that a large set of leading anomalous dimensions is determined up to one parameter, in contrast to the $n=1$ case where the anomalous dimensions are fixed. Like the $\phi^{5}$ theory, the $\phi^{2n+1}$ theories with higher $n$ are also associated with multicritical Yang-Lee models vonGehlen:1994rp ; Lencses:2022ira ; Lencses:2023evr ; Lencses:2024wib , and they provide a series of universality classes containing multicritical points with $\mathbb{Z}_{2}$-odd interactions. See also Zambelli:2016cbw ; Gracey:2017okb ; Codello:2017qek ; Codello:2017epp for related studies of the canonical case $k=1$. This concludes our discussion of the theories that are investigated in this paper.
@@
where the indices are suppressed for simplicity. The partially conserved currents do not vanish when we act with fewer derivatives. These operators are related to the higher-spin symmetry of the free theory. One can construct conserved charges from the partially conserved currents, generalizing the $k=1$ higher-spin algebra. We refer to Brust:2016gjy and references therein for more details. For $k>1$, the interacting versions of the partially conserved currents play a significant role in the analytic bootstrap, as their twists are lower than those of the usual conserved currents. In the lightcone limit, the interacting counterparts of the partially conserved currents are more important than the Regge trajectories associated with the conserved currents.
@@
To derive the anomalous dimensions of the bilinear operators, we employ the conformal multiplet recombination method initiated in Rychkov:2015naa , which considered the canonical $\phi^{4}$ theory. In Rychkov:2015naa , it was further suggested that the results can be extended to the canonical $\phi^{3}$ theory at $d=6-\epsilon$.  [^12] According to Rychkov:2015naa ; Gliozzi:2016ysv ; Gliozzi:2017gzh , Yu Nakayama has studied the $\epsilon$ expansion using the multiplet recombination in his unpublished notes, such as the $\epsilon=6-d$ expansion of the $\phi^{3}$ theory. A key difference is the absence of $\mathbb{Z}_{2}$ symmetry in the $\phi^{3}$ theory, in contrast to the $\phi^{4}$ case. Since the procedure in Rychkov:2015naa relies on the $\mathbb{Z}_{2}$ symmetry, its generalization to the $\mathbb{Z}_{2}$-odd cases does not seem straightforward. A similar approach based on the equation of motion was proposed to study the canonical $\phi^{3}$ theory Nii:2016lpa . By applying the equation of motion to correlation functions, the constraints on the anomalous dimensions can be derived more readily. In this paper, we revisit the $\phi^{2n+1}$ theory from the multiplet recombination perspective and extend the results to some spinning operators. As a consistency check, we also verify that the results for the anomalous dimensions and operator product expansion (OPE) coefficients are consistent with crossing symmetry using the analytic bootstrap.
@@
The multiplet recombination asserts that free multiplets recombine into long multiplets at interacting IR fixed points. For instance, at $d=4-\epsilon$, the Wilson-Fisher multiplet $\{\phi\}$ is a long multiplet, into which the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{3}_{\text{f}}\}$ recombine. Here the subscript f implies that the operators are associated with the free theory. In the Gaussian limit, the descendant $\Box\phi$ is expected to become null and decouple, but a singular change in the normalization can lead to a physical operator, which is identified with the free-theory primary operator $\phi^{3}_{\text{f}}$. In other words, the Wilson-Fisher counterpart of $\phi^{3}_{\text{f}}$ is identified as a descendant $\Box\phi$ up to a normalization factor  [^13] All physical operators in the free theory should have counterparts in the interacting theory.:
@@
Despite the resemblance to the equation of motion $\Box\phi\propto\epsilon\,\phi^{3}$, the multiplet recombination is conceptually different, as it can be formulated purely in the CFT language and does not require a Lagrangian description. The multiplet recombination above generalizes to various cases. As in Guo:2023qtt , we follow a suggestion of Gliozzi:2017gzh and apply the multiplet recombination method to the $\mathbb{Z}_{2}$-odd deformations of the free higher-derivative CFTs. See Basu:2015gpa ; Yamaguchi:2016pbj ; Ghosh:2015opa ; Raju:2015fza ; Gliozzi:2016ysv ; Roumpedakis:2016qcg ; Soderberg:2017oaa ; Behan:2017emf ; Gliozzi:2017hni ; Gliozzi:2017gzh ; Nishioka:2022odm ; Nishioka:2022qmj ; Antunes:2022vtb ; Guo:2023qtt for applications of the multiplet recombination to numerous CFTs, and Safari:2017irw ; Nii:2016lpa ; Hasegawa:2016piv ; Hasegawa:2018yqg ; Skvortsov:2015pea ; Giombi:2016hkj ; Giombi:2017rhm ; Giombi:2016zwa ; Codello:2017qek ; Codello:2018nbe ; Antipin:2019vdg ; Vacca:2019rsh ; Giombi:2020rmc ; Giombi:2020xah ; Dey:2020jlc ; Giombi:2021cnr ; Safari:2021ocb ; Zhou:2022pah ; Bissi:2022bgu ; Herzog:2022jlx ; Giombi:2022vnz ; SoderbergRousu:2023pbe for applications with more emphasis on the equations of motion.
This paper is organized as follows. In section 2, we consider the generalized Yang-Lee edge singularity. We first employ the multiplet recombination to compute the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}$, and then we use the analytic bootstrap to verify that the results are consistent with crossing symmetry. In section 3 we carry out a similar analysis, obtaining the anomalous dimensions in the generalized Potts model, and we check that the results are consistent with crossing symmetry as well. Section 4 focuses on the generalized $\phi^{2n+1}$ theory with $n>1$, where we only consider the single-field case. We determine the anomalous dimensions up to a free parameter, and then we verify the consistency between the multiplet-recombination results and crossing symmetry. We summarize the main results in table 1. In appendix A we work out the action of Laplacians on the scalar three-point function. In appendix B we verify that the multiplet-recombination results in section 2, 3, and 4 are consistent with crossing symmetry at subleading twist. We briefly review some representation theory of the symmetric group in appendix C. More details about the tensor structures in the Potts model can be found in appendix D. In appendix E we compute the ratios of OPE coefficients relevant to the multiplet recombination. In appendix F we derive an OPE coefficient used in the analytic bootstrap. In appendix G we discuss the $N\rightarrow 0$ limit of the generalized Potts model in more detail, especially the logarithmic structures. Appendix H discusses some special limits in the generalized O($N$)-symmetric $\phi^{2n}$ theory.
@@ 2 Generalized Yang-Lee edge singularity
Let us first consider the generalized Yang-Lee CFT. At $d=d_{\text{u}}-\epsilon$, the action reads [^14] We do not need to consider derivative interactions. The reason is that a derivative interaction should be constructed by fewer $\phi$’s, namely, two $\phi$’s, but a derivative term with two $\phi$’s is just the kinetic term, up to a total derivative.
$S\propto\int\mathrm{d}^{d}x\left(\phi\,\Box^{k}\phi+g\mu^{\epsilon/2}\phi^{3}\right)\,,$ (5)
@@ 2.1 Multiplet recombination
where $\alpha=\alpha(\epsilon)$ vanishes in the Gaussian limit, i.e., $\lim_{\epsilon\rightarrow 0}\alpha=0$. The above equation means that we identify the descendant $\alpha^{-1}\Box^{k}\phi$ as the deformed version of $\phi^{2}_{\text{f}}$, and the normalization $\alpha^{-1}$ is introduced to ensure a finite Gaussian limit. In other words, the free multiplets $\{\phi_{\text{f}}\}$ and $\{\phi^{2}_{\text{f}}\}$ recombine into a long multiplet $\{\phi\}$.
@@
Our focus will be on the two-point and three-point functions. We study the matching conditions for the scalar correlators first. As we will see below, these conditions determine the anomalous dimension $\gamma_{\phi}$ and the normalization $\alpha$, as well as the OPE coefficient associated with $\phi\in\phi\times\phi$, at leading order in the $\epsilon$ expansion. Then, we consider the matching condition for the spinning correlator involving a primary bilinear operator. [^15] The spin of the bilinear operator can be zero. Using the results for $\gamma_{\phi}$ and $\alpha$, we obtain the leading anomalous dimension of the bilinear operator.
@@
On the right-hand side of (7), the free correlator is given by Wick contractions.  [^16] We assume that the Gaussian limit is smooth, so the normalization in the free theory is also fixed by (9), i.e., $\langle\phi_{\text{f}}(x_{1})\phi_{\text{f}}(x_{2})\rangle=1/x_{12}^{\smash{2\Delta_{\phi_{\text{f}}}}}$. The functional forms in $x_{12}$ indeed match in (8), since
@@
and substitute $d=6k-\epsilon$ into (8). Omitting the functional forms in $x_{12}$, we are left with
@@
As before, the right-hand side is computed using Wick contractions. On the left-hand side, we act on $\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ with the Laplacians. The result is worked out in appendix A. We obtain
@@
where $\lambda_{\phi\phi\phi}$ is the OPE coefficient of $\phi\in\phi\times\phi$. The coefficients $B_{i,j}$ and $C_{i,j}^{a,b}$ are given in (111) and (112). Due to the last Pochhammer symbol in (112), $C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}$ is proportional to the anomalous dimension $\gamma_{\phi}$, except for the $i=j=k$ case where $C_{i,j}^{\Delta_{\smash{\phi}},\Delta_{\smash{\phi}}}$ is finite in the Gaussian limit. In other words, the dominant term in the $\epsilon\rightarrow 0$ limit is given by $i=j=k$. Then, one can check that the functional forms match in (15). We obtain a constraint on $\lambda_{\phi\phi\phi}$ and $\alpha$,
@@
indicating that the OPE coefficient $\lambda_{\phi\phi\phi}$ is of order $\alpha$. So far, we have obtained two constraints, i.e., (13) and (16), but they involve three unknown parameters: $\alpha$, $\gamma_{\phi}$, and $\lambda_{\phi\phi\phi}$. We would like to deduce another constraint that does not involve additional unknown parameters, so we consider the three-point function $\langle\phi(x_{1})\phi(x_{2})\phi(x_{3})\rangle$ again. In (14), we have considered the matching conditions associated with one type of $\Box^{k}$. It is natural to introduce more $\Box^{k}$’s. However, as a result of the $\mathbb{Z}_{2}$ symmetry of the free theory, the free correlator $\langle\phi^{2}_{\text{f}}(x_{1})\phi^{2}_{\text{f}}(x_{2})\phi_{\text{f}}(x_{3})\rangle$ vanishes and thus the matching condition associated with $\langle\Box^{k}\phi(x_{1})\Box^{k}\phi(x_{2})\phi(x_{3})\rangle$ is not particularly useful. If we consider three types of $\Box^{k}$’s, the corresponding free-theory correlator respects the $\mathbb{Z}_{2}$ symmetry, so this is a useful matching condition,
@@
The action of the Laplacians is derived in appendix A. The explicit result associated with three types of $\Box^{k}$’s can be found in (A). As before, the functional forms match, leaving us with a constraint on $\gamma_{\phi}$, $\lambda_{\phi\phi\phi}$, and $\alpha$: